# Transport, Variational Inference and Diffusions: with Applications to Annealed Flows and Schrödinger Bridges

### v1

@@ Abstract
This paper explores the connections between optimal transport and variational inference, with a focus on forward and reverse time stochastic differential equations and Girsanov transformations.We present a principled and systematic framework for sampling and generative modelling centred around divergences on path space. Our work culminates in the development of a novel score-based annealed flow technique (with connections to Jarzynski and Crooks identities from statistical physics) and a regularised iterative proportional fitting (IPF)-type objective, departing from the sequential nature of standard IPF. Through a series of generative modelling examples and a double-well-based rare event task, we showcase the potential of the proposed methods.
Machine Learning, ICML
[…]
@@ 1 Introduction
Optimal transport (Villani et al., 2009) and variational inference (Blei et al., 2017) have for a long time been separate fields of research. In recent years, many fruitful connections have been established (Liu et al., 2019), in particular based on dynamical formulations (Tzen and Raginsky, 2019), and in conjunction with time reversals (Huang et al., 2021a; Song et al., 2021). In this paper, we enhance those relationships based on forward and reverse time stochastic differential equations, and associated Girsanov transformations.
To set the stage, we recall a classical approach (Kingma and Welling, 2014; Rezende and Mohamed, 2015) towards generating samples from a target distribution $\mu({\bm{x}})$, which is the goal both in generative modelling and sampling:
@@
${\bm{x}}\sim\mu({\bm{x}}),\quad{\bm{z}}|{\bm{x}}\sim q^{\phi}({\bm{z}}|{\bm{x}}),$ (3)
relying on an appropriately parameterised backward transition $q^{\phi}({\bm{z}}|{\bm{x}})$. We will say that (1) and (3) are reversals of each other in the case when the joint distributions associated to both coincide, that is, when
$q^{\phi}({\bm{z}}|{\bm{x}})\mu({\bm{x}})=p^{\theta}({\bm{x}}|{\bm{z}})\nu({\bm{z}}).$ (4)
To appreciate the significance of the reversed process in (3), notice that if the reversal relation in (4) holds, then (2) is implied by integrating both sides with respect to ${\bm{z}}$. Building on this observation, it is natural to define the loss function
@@
where $D$ is a divergence [^1] As usual, divergences are characterised by the requirement that $D(\alpha\big{|}\big{|}\beta)\geq 0$, with equality if and only if $\alpha=\beta$. between distributions yet to be specified. Along the lines of Bengio et al. (2021); Sohl-Dickstein et al. (2015); Wu et al. (2020); Liu et al. (a), we summarise the discussion so far as follows, laying the foundation for algorithmic approaches that aim at sampling from $\mu({\bm{x}})$ by minimising $\mathcal{L}_{D}(\phi,\theta)$:
@@
$\mu({\bm{x}})\!=\!\!\int\!\!p^{\theta}({\bm{x}}|{\bm{z}})\nu(\mathrm{d}{\bm{z}})\;\text{and}\;\nu({\bm{z}})\!\!=\!\!\int\!\!q^{\phi}({\bm{z}}|{\bm{x}})\mu(\mathrm{d}{\bm{x}}),$ (6)
@@
KL-divergence, ELBO and variational inference. Choosing $D=D_{\mathrm{KL}}$ in (5), variational inference (VI) and latent variable model (LVM) based approaches (Dempster et al., 1977; Blei et al., 2017; Kingma and Welling, 2014) can elegantly be placed within Framework 1. Indeed, direct computation shows that
$\mathcal{L}_{D_{\mathrm{KL}}}(\phi,\theta)=-$ $\mathbb{E}_{{\bm{x}}\sim\mu({\bm{x}})}\overbrace{\left[\int\ln\frac{p^{\theta}({\bm{x}}|{\bm{z}})\nu({\bm{z}})}{q^{\phi}({\bm{z}}|{\bm{x}})}q^{\phi}(\mathrm{d}{\bm{z}}|{\bm{x}})\right]}^{=\mathrm{ELBO}_{x}(\phi,\theta)}$
$+\int\operatorname{ln}\mu({\bm{x}})\mu(\mathrm{d}{\bm{x}}),$ (7)
@@
so that minimising $\mathcal{L}_{D_{\mathrm{KL}}}(\phi,\theta)$ is equivalent to maximising the expected evidence lower bound (ELBO), also known as the negative free energy (Blei et al., 2017). This derivation is alternative to the standard approach via maximum likelihood and convex duality (or Jensen’s inequality) (Kingma et al., 2021, Section 2.2), and straightforwardly accomodates various modifications by replacing the $D_{\mathrm{KL}}$-divergence (see Appendix B). Other popular choices for $D$ include the $f$ and $\alpha$ divergences (Csiszár, 1967; Ali and Silvey, 1966; Rényi, 1961). Couplings, (optimal) transport and nonuniqueness. Assuming that (4) holds, it is natural to define the joint distribution $\pi({\bm{x}},{\bm{z}}):=q^{\phi}({\bm{z}}|{\bm{x}})\mu({\bm{x}})=p^{\theta}({\bm{x}}|{\bm{z}})\nu({\bm{z}})$, which is then a coupling between $\mu({\bm{x}})$ and $\nu({\bm{z}})$ by virtue of (6). Viewed from this angle, it is clear that the set of minimisers associated to $\mathcal{L}(\phi,\theta)$ stands in one-to-one correspondence with the set of couplings between $\mu({\bm{x}})$ and $\nu({\bm{z}})$, provided that the parameterisations are chosen flexibly enough. Under the latter assumption, the objective in (5) hence generically admits an infinite number of minimisers, rendering algorithmic approaches solely based on Framework 1 potentially unstable and their output hard to interpret. In the language of optimal transport (Villani, 2003), minimising $\mathcal{L}(\phi,\theta)$ enforces the marginal (‘transport’) constraints in (6) without a selection principle based on an appropriate cost function (‘optimal’).
@@
In practice, methodologies such as VAEs (Kingma and Welling, 2014) parameterise the transition densities $p^{\theta}({\bm{x}}|{\bm{z}})$ and $q^{\phi}({\bm{z}}|{\bm{x}})$ with a restricted family of distributions (such as Gaussians), thus restricting the set of couplings. Expectation maximisation (EM) carries out the minimisation of $\mathcal{L}(\phi,\theta)$ in a component-wise fashion, resolving nonquniqueness in a procedural manner (see Section 4.1 for a connection to Schrödinger bridges). In this paper, we develop principled approaches towards resolving the nonuniqueness inherent in Framework 1, relying on a hierarchical extension of $p^{\theta}({\bm{x}}|{\bm{z}})$ and $q^{\phi}({\bm{z}}|{\bm{x}})$ and proceeding to an infinite-depth limit. In particular, we proceed as follows:
@@
Outline and contributions. In Section 2 we recall hierarchical VAEs (Rezende et al., 2014) and, following Tzen and Raginsky (2019), proceed to the infinite-depth limit described by the SDEs in (2.2). Readers more familiar with VI and discrete time might want to take the development in Section 2.1 as an explanation of (2.2); readers with background in stochastic analysis might take Framework ${}^{\prime}$ as their starting point. In Proposition 2.2 we provide a generalised form of the Girsanov theorem for forward-reverse time SDEs, crucially incorporating the choice of a reference process that allows us to reason about sampling and generation in a systematic and principled way. In Section 3, we utilise the flexibility in this formulation to recover recent approaches from the literature, allowing for slight variations and generalisations, and develop a novel score-based annealed flow technique. Section 4 reinforces the connection between optimal transport and VI by establishing a correspondence between expectation-maximisation (EM) and iterative proportional fitting (IPF). Motivated by this result and building on the set-up of Section 2, we suggest a regularised IPF-type objective that allows us to circumvent the sequential nature of standard IPF. Finally, we explore our proposed regularised IPF objective across a series of toy generative modelling examples and a double-well-based rare event (Hartmann et al., 2013) task.
2 From hierarchical VAEs to forward-reverse SDEs
@@
$q^{\phi}({\bm{y}}_{L},{\bm{y}}_{L-1},\ldots,{\bm{y}}_{1}|{\bm{y}}_{0}):=\prod_{l=1}^{L}q^{\phi_{l-1}}({\bm{y}}_{l}|{\bm{y}}_{l-1}),$
$p^{\theta}({\bm{y}}_{0},{\bm{y}}_{1},\ldots,{\bm{y}}_{L-1}|{\bm{y}}_{L}):=\prod_{l=1}^{L}p^{\theta_{l}}({\bm{y}}_{l-1}|{\bm{y}}_{l}),$ (8)
@@
so that $q^{\phi}({\bm{z}}|{\bm{x}})$ and $p^{\theta}({\bm{x}}|{\bm{z}})$ can be obtained from (2.1) by marginalising over the auxiliary variables ${\bm{y}}_{1},\ldots,{\bm{y}}_{L-1}$. Here, $\phi=(\phi_{0},\ldots,\phi_{L-1})$ and $\theta=(\theta_{1},\ldots,\theta_{L})$ refer to sets of parameters to be specified in more detail below. Further introducing notation, we write $q^{\mu,\phi}({\bm{y}}_{0:L}):=q^{\phi}({\bm{y}}_{1:L}|{\bm{y}}_{0})\mu({\bm{y}}_{0})$ as well as $p^{\nu,\theta}({\bm{y}}_{0:L}):=p^{\theta}({\bm{y}}_{0:L-1}|{\bm{y}}_{L})\nu({\bm{y}}_{L})$ and think of those implied joint distributions as emanating from $\mu({\bm{x}})=\mu({\bm{y}}_{0})$ and $\nu({\bm{z}})=\nu({\bm{y}}_{L})$, respectively, moving ‘forwards’ or ‘backwards’ according to the specific choices for $\phi$ and $\theta$. In the regime when $L$ is large, the models in (2.1) are very expressive, even if the intermediate transition kernels are parameterised in a simple manner. We hence proceed by assuming Gaussian distributions,
$q^{\phi_{l-1}}({\bm{y}}_{l}|{\bm{y}}_{l-1})={\mathcal{N}}({\bm{y}}_{l}|{\bm{y}}_{l-1}+\delta a^{\phi}_{l-1}({\bm{y}}_{l-1}),\delta\sigma^{2}I),$
$p^{\theta_{l}}({\bm{y}}_{l-1}|{\bm{y}}_{l})={\mathcal{N}}({\bm{y}}_{l-1}|{\bm{y}}_{l}+\delta b^{\theta}_{l}({\bm{y}}_{l}),\delta\sigma^{2}I),$ (9)
@@
where $\sigma>0$ controls the standard deviation, and $\delta>0$ is a small parameter, anticipating the limits $L\rightarrow\infty$, $\delta\rightarrow 0$ to be taken in Section 2.2 below. The vector fields $a^{\phi}_{l}({\bm{y}}_{l})$ and $b_{l}^{\theta}({\bm{y}}_{l})$ introduced in (2.1) should be thought of as parameterised by $\phi$ and $\theta$, but we will henceforth suppress this dependence for notational efficiency (in fact, those vector fields form a natural parameterisation in their own right).
The models (2.1)-(2.1) could equivalently be defined via the Markov chains
${\bm{y}}_{l+1}$ $={\bm{y}}_{l}+\delta a_{l}({\bm{y}}_{l})+\sqrt{\delta}\sigma\xi_{l},\quad{\bm{y}}_{0}\sim\mu$
$\implies{\bm{y}}_{0:L}\sim q^{\mu,\phi}({\bm{y}}_{0:L}),$ (10a)
${\bm{y}}_{l-1}$ $={\bm{y}}_{l}+\delta b_{l}({\bm{y}}_{l})+\sqrt{\delta}\sigma\xi_{l},\quad{\bm{y}}_{L}\sim\nu$
$\implies{\bm{y}}_{0:L}\sim p^{\nu,\theta}({\bm{y}}_{0:L}),$ (10b)
@@
where $(\xi_{l})_{l=1}^{L}$ is an iid sequence of standard Gaussian random variables. As indicated, the forward process in (2.1) may serve to define the distribution $q^{\mu,\phi}({\bm{y}}_{0:L})$, whilst the backward process in (2.1) induces $p^{\nu,\theta}({\bm{y}}_{0:L})$. Note that the transition densities $p^{\theta}({\bm{x}}|{\bm{z}})$ and $q^{\phi}({\bm{z}}|{\bm{x}})$ obtained as the marginals of (2.1) will in general not be available in closed form. However, generalising slightly from Framework 1, we may set out to minimise the extended loss
@@
$D_{\mathrm{KL}}($ $q^{\mu,\phi}({\bm{y}}_{0:L})||p^{\nu,\theta}({\bm{y}}_{0:L}))\geq$
$D_{\mathrm{KL}}\left(q^{\phi}({\bm{z}}|{\bm{x}})\mu({\bm{x}})\big{|}\big{|}p^{\theta}({\bm{x}}|{\bm{z}})\nu({\bm{z}})\right),$ (12)
@@
2.2 SDE transition densities – hierarchical VAEs in the infinite depth limit

### v6

# Transport meets Variational Inference: Controlled Monte Carlo Diffusions

@@ Abstract
Connecting optimal transport and variational inference, we present a principled and systematic framework for sampling and generative modelling centred around divergences on path space. Our work culminates in the development of Controlled Monte Carlo Diffusions for sampling and inference, a score-based annealing technique that crucially adapts both forward and backward dynamics in a diffusion model. On the way, we clarify the relationship between the EM-algorithm and iterative proportional fitting (IPF) for Schrödinger bridges, providing a conceptual link between fields. Finally, we show that CMCD has a strong foundation in the Jarzinsky and Crooks identities from statistical physics, and that it convincingly outperforms competing approaches across a wide array of experiments.
{NoHyper}\$*\$\$*\$footnotetext: Equal contribution.
@@ 1 Introduction
Optimal transport (Villani et al., 2009) and variational inference (Blei et al., 2017) have for a long time been separate fields of research. In recent years, many fruitful connections have been established (Liu et al., 2019), in particular based on dynamical formulations (Tzen & Raginsky, 2019a), and in conjunction with time reversals (Huang et al., 2021a; Song et al., 2021). The goal of this paper is twofold: In the first part, we enhance those relationships based on forward and reverse time diffusions, and associated Girsanov transformations, arriving at a unifying framework for generative modeling and sampling. In the second part, we build on this and develop a novel score-based scheme for sampling from unnormalised densities. To set the stage, we recall a classical approach (Kingma & Welling, 2014; Rezende & Mohamed, 2015) towards generating samples from a target distribution $\mu({\bm{x}})$, which is the goal both in generative modelling and sampling:
@@
${\bm{x}}\sim\mu({\bm{x}}),\qquad{\bm{z}}|{\bm{x}}\sim q^{\phi}({\bm{z}}|{\bm{x}}),$ (3)
relying on an appropriately parameterised backward transition $q^{\phi}({\bm{z}}|{\bm{x}})$. We will say that (1) and (3) are reversals of each other in the case when their joint distributions coincide, that is, when
$q^{\phi}({\bm{z}}|{\bm{x}})\mu({\bm{x}})=p^{\theta}({\bm{x}}|{\bm{z}})\nu({\bm{z}}).\vspace{-0.1cm}$ (4)
To appreciate the significance of (3), notice that if (4) holds, then (2) is implied by integrating both sides with respect to ${\bm{z}}$. Building on this observation, it is natural to define the loss function
@@
where $D$ is a divergence [^1] As usual, divergences are characterised by the requirement that $D(\alpha\big{|}\big{|}\beta)\geq 0$, with equality iff $\alpha=\beta$. between distributions yet to be specified. Along the lines of Bengio et al. (2021); Sohl-Dickstein et al. (2015); Wu et al. (2020); Liu et al. (b), we have now laid the foundations for algorithmic approaches that aim at sampling from $\mu({\bm{x}})$ by minimising $\mathcal{L}_{D}(\phi,\theta)$:
@@
$\mu({\bm{x}})\!=\!\!\int\!\!p^{\theta}({\bm{x}}|{\bm{z}})\nu(\mathrm{d}{\bm{z}})\;\quad\text{and}\quad\;\nu({\bm{z}})\!\!=\!\!\int\!\!q^{\phi}({\bm{z}}|{\bm{x}})\mu(\mathrm{d}{\bm{x}}),$ (6)
@@
The sampling problem. Let $\nu$ denote a probability density function on ${\mathbb{R}}^{d}$ of the form ${\nu({\bm{z}})=\frac{\hat{\nu}({\bm{z}})}{Z},\quad Z=\int_{\mathbb{R}^{d}}\hat{\nu}({\bm{z}})\mathrm{d}{\bm{z}},}$ where $\hat{\nu}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{+}$can be differentiated and evaluated pointwise but the normalizing constant $Z$ is intractable. We are interested in both estimating $Z$ and obtaining approximate samples from $\nu$ given we can sample from a more tractable density $\mu$. Framework 1 provides us with an objective to tackle the sampling problem as once $\mathcal{L}_{D}(\phi,\theta)=0$, we can generate samples from $\nu({\bm{z}})$ via the variational distribution $q^{\phi}({\bm{z}}|{\bm{x}})$. Through variational inference and optimal transport, we discuss relationships to classical methods as well as shortcomings:
@@
KL-divergence, ELBO and variational inference. Choosing $D=D_{\mathrm{KL}}$ in (5), variational inference (VI) and latent variable model based approaches (Dempster et al., 1977; Blei et al., 2017; Kingma & Welling, 2014) can elegantly be placed within Framework 1. Indeed, direct computation (see Appendix B) shows that $\mathcal{L}_{D_{\mathrm{KL}}}(\phi,\theta)=-\mathbb{E}_{{\bm{x}}\sim\mu({\bm{x}})}[\mathrm{ELBO}_{x}(\phi,\theta)]+\mathbb{E}_{{\bm{x}}\sim\mu({\bm{x}})}[\operatorname{ln}\mu(x)]$, so that minimising $\mathcal{L}_{D_{\mathrm{KL}}}(\phi,\theta)$ is equivalent to maximising the expected evidence lower bound (ELBO), also known as the negative free energy (Blei et al., 2017). This derivation is alternative to the standard approach via maximum likelihood and convex duality (or Jensen’s inequality) (Kingma et al., 2021, Section 2.2), and directly accomodates various modifications by replacing the $D_{\mathrm{KL}}$-divergence (see Appendix B).
@@
Couplings, (optimal) transport and nonuniqueness. Assuming (4) holds, it is natural to define the joint distribution $\pi({\bm{x}},{\bm{z}}):=q^{\phi}({\bm{z}}|{\bm{x}})\mu({\bm{x}})=p^{\theta}({\bm{x}}|{\bm{z}})\nu({\bm{z}})$, which is a coupling between $\mu({\bm{x}})$ and $\nu({\bm{z}})$. Viewed from this angle, the set of minimisers of $\mathcal{L}(\phi,\theta)$ stands in one-to-one correspondence with the set of couplings between $\mu({\bm{x}})$ and $\nu({\bm{z}})$, provided that the parameterisations are chosen flexibly enough. Under the latter assumption, the objective in (5) admits an infinite number of minimisers, rendering algorithmic approaches solely based on Framework 1 potentially unstable and their output hard to interpret. In the language of optimal transport (Villani, 2003), minimising $\mathcal{L}(\phi,\theta)$ enforces the marginal (‘transport’) constraints in (6) without a selection principle based on an appropriate cost function (‘optimal’).
@@
Methods such as VAEs (Kingma & Welling, 2014) parameterise $p^{\theta}({\bm{x}}|{\bm{z}})$ and $q^{\phi}({\bm{z}}|{\bm{x}})$ with a restricted family of distributions (such as Gaussians), thus restricting the set of couplings. Expectation maximisation (EM) minimises $\mathcal{L}(\phi,\theta)$ in a component-wise fashion, resolving nonquniqueness in a procedural manner (see Section 3.1). Common diffusion models fix either $p^{\theta}({\bm{x}}|{\bm{z}})$ or $q^{\phi}({\bm{z}}|{\bm{x}})$, and thus select a coupling (Section 2.2). In this paper, we argue that the full potential of diffusion models can be unleashed by training the forward and backward processes at the same time, but appropriate modifications that resolve the nonuniqueness inherent in Framework 1 need to be imposed. To develop principled approaches towards this, we proceed as follows:
@@
Outline and contributions. In Section 2 we recall hierarchical VAEs (Rezende et al., 2014) and, following Tzen & Raginsky (2019a), proceed to the infinite-depth limit described by the SDEs in (12). Readers more familiar with VI and discrete time might want to take the development in Section 2.1 as an explanation of (12); readers with background in stochastic analysis might take Framework 1 as their starting point. In Proposition 2.2 we provide a generalised form of the Girsanov theorem for forward-reverse time SDEs, crucially incorporating the choice of a reference process that allows us to reason about sampling and generation in a systematic and principled way. We demonstrate that a range of widely used approaches, such as score-based diffusions and path integral samplers, among others, are special cases of our unifying framework (Section 2.2). Similarly in Section 3.1 we unify optimal transport (OT) and VI under our framework by establishing a correspondence between expectation-maximisation (EM) and iterative proportional fitting (IPF). Going further, we show that this framework allows us to derive new methods:
In Section 3.2, we derive a novel score-based annealed flow technique, the Controlled Monte Carlo Diffusion (CMCD) sampler, and show that it may be viewed as an infinitesimal analogue of the method from Section 3.1. Finally, we connect CMCD to the foundational identities by Crooks and Jarzynki in statistical physics, and show that it empirically outperforms a range of state-of-the-art inference methods in sampling and estimating normalizing constants (Section 4).
2 From hierarchical VAEs to forward-reverse time diffusions
@@
$q^{\phi}({\bm{y}}_{L},\ldots,{\bm{y}}_{1}|{\bm{y}}_{0}):=\prod_{l=1}^{L}q^{\phi_{l-1}}({\bm{y}}_{l}|{\bm{y}}_{l-1}),\qquad p^{\theta}({\bm{y}}_{0},\ldots,{\bm{y}}_{L-1}|{\bm{y}}_{L}):=\prod_{l=1}^{L}p^{\theta_{l}}({\bm{y}}_{l-1}|{\bm{y}}_{l}),\vspace{-0.2cm}$ (7)
@@
so that $q^{\phi}({\bm{z}}|{\bm{x}})$ and $p^{\theta}({\bm{x}}|{\bm{z}})$ can be obtained from (7) by marginalising over the auxiliary variables ${\bm{y}}_{1},\ldots,{\bm{y}}_{L-1}$. Here, $\phi=(\phi_{0},\ldots,\phi_{L-1})$ and $\theta=(\theta_{1},\ldots,\theta_{L})$ refer to sets of parameters to be specified in more detail below. Further introducing notation, we write $q^{\mu,\phi}({\bm{y}}_{0:L}):=q^{\phi}({\bm{y}}_{1:L}|{\bm{y}}_{0})\mu({\bm{y}}_{0})$ as well as $p^{\nu,\theta}({\bm{y}}_{0:L}):=p^{\theta}({\bm{y}}_{0:L-1}|{\bm{y}}_{L})\nu({\bm{y}}_{L})$ and think of those implied joint distributions as emanating from $\mu({\bm{x}})=\mu({\bm{y}}_{0})$ and $\nu({\bm{z}})=\nu({\bm{y}}_{L})$, respectively, moving ‘forwards’ or ‘backwards’ according to the specific choices for $\phi$ and $\theta$. In the regime when $L$ is large, the models in (7) are very expressive, even if the intermediate transition kernels are parameterised in a simple manner. We hence proceed by assuming Gaussian distributions,
$q^{\phi_{l-1}}\!({\bm{y}}_{l}|{\bm{y}}_{l-1}\!)\!=\!{\mathcal{N}}({\bm{y}}_{l}|{\bm{y}}_{l-1}\!\!+\!\delta a^{\phi}_{l-1}({\bm{y}}_{l-1}),\delta\sigma^{2}\!I),\;\;p^{\theta_{l}}\!({\bm{y}}_{l-1}|{\bm{y}}_{l}\!)\!=\!{\mathcal{N}}({\bm{y}}_{l-1}|{\bm{y}}_{l}\!\!+\!\delta b^{\theta}_{l}({\bm{y}}_{l}),\delta\sigma^{2}\!I),$ (8)
@@
where $\sigma>0$ controls the standard deviation, and $\delta>0$ is a small parameter, anticipating the limits $L\rightarrow\infty$, $\delta\rightarrow 0$ to be taken in Section 2.2 below. The vector fields $a^{\phi}_{l}({\bm{y}}_{l})$ and $b_{l}^{\theta}({\bm{y}}_{l})$ introduced in (8) should be thought of as parameterised by $\phi$ and $\theta$, but we will henceforth suppress this for brevity.
The models (7)-(8) could equivalently be defined via the Markov chains
${\bm{y}}_{l+1}$ $={\bm{y}}_{l}+\delta a_{l}({\bm{y}}_{l})+\sqrt{\delta}\sigma\xi_{l},\qquad{\bm{y}}_{0}\sim\mu\implies{\bm{y}}_{0:L}\sim q^{\mu,\phi}({\bm{y}}_{0:L}),$ (9a)
${\bm{y}}_{l-1}={\bm{y}}_{l}+\delta b_{l}({\bm{y}}_{l})+\sqrt{\delta}\sigma\xi_{l},\qquad{\bm{y}}_{L}\sim\nu\implies{\bm{y}}_{0:L}\sim p^{\nu,\theta}({\bm{y}}_{0:L}),$ (9b)
@@
where $(\xi_{l})_{l=1}^{L}$ is an iid sequence of standard Gaussian random variables. As indicated, the forward process in (9a) may serve to define the distribution $q^{\mu,\phi}({\bm{y}}_{0:L})$, whilst the backward process in (9b) induces $p^{\nu,\theta}({\bm{y}}_{0:L})$. Note that the transition densities $p^{\theta}({\bm{x}}|{\bm{z}})$ and $q^{\phi}({\bm{z}}|{\bm{x}})$ obtained as the marginals of (7) will in general not be available in closed form. However, generalising slightly from Framework 1, we may set out to minimise the extended loss
@@
$D_{\mathrm{KL}}(q^{\mu,\phi}({\bm{y}}_{0:L})||p^{\nu,\theta}({\bm{y}}_{0:L}))\geq D_{\mathrm{KL}}\left(q^{\phi}({\bm{z}}|{\bm{x}})\mu({\bm{x}})\big{|}\big{|}p^{\theta}({\bm{x}}|{\bm{z}})\nu({\bm{z}})\right),$ (11)
@@
2.2 Diffusion models – hierarchical VAEs in the infinite depth limit